# Novel method for in-situ drift velocity measurement in large volume TPCs: the Geometry Reference Chamber of the NA61/SHINE experiment at CERN

### v1

@@ 4 Drift velocity calibration using the GRC
Here $y_{\mathrm{nom}}$ is the reconstructed drift coordinate assuming a nominal estimate during reconstruction for the position of the TPC amplification plane at $y_{\mathrm{anode},\mathrm{nom}}$. The symbol $t_{0,\mathrm{nom}}$ stands for a nominal estimate for the trigger latency during reconstruction. $t_{\mathrm{drift,raw}}$ is the measured particle track cluster position in terms of drift time as measured by the TPC readout. $v_{\mathrm{drift,nom}}$ is the nominal estimate for the electron drift velocity in the TPC during reconstruction. The quantities labeled by $\,{()}_{\mathrm{true}}\,$ are the true but unknown values of these calibration factors. The negative sign in Eq.(4.2) is due to the coordinate convention in NA61/SHINE. The coordinate $y_{\mathrm{nom}}$ measured by the TPC along the drift direction corresponds to a GRC measurement $y_{\mathrm{true}}$, for each track matched to a GRC hit, as seen in Fig.1. The matching is done via a typically ${\sim}3\,\mathrm{cm}$ wide tolerance window. By eliminating the running variable $t_{\mathrm{drift,raw}}$ in Eq.(4.2), one infers
@@
for the ensemble of TPC-GRC tracks. Here $\Delta{y}:=y_{\mathrm{nom}}-y_{\mathrm{true}}$ stands for the drift coordinate TPC-GRC mismatch and $y_{\mathrm{true}}$ stands for the GRC measurement. Occasionally, the label $()_{\mathrm{true}}$ will be suppressed where not confusing. The slope of the $\Delta{y}$ versus $y$ scatter plot determines the correction factor between the true drift velocity $v_{\mathrm{drift,true}}$ and the estimate $v_{\mathrm{drift,nom}}$ assumed during the reconstruction, thus $v_{\mathrm{drift,true}}$ can be determined from the TPC-GRC mismatch data. This is shown in Fig.5 top panel for a typical low-multiplicity data set (proton-carbon collisions at $120\,\mathrm{GeV}/c$ beam momentum, recorded in 2023). In Fig.5 bottom left panel the same is shown, but for a typical high-multiplicity data set (minimum-bias Pb+Pb collisions at $150\,A\mathrm{GeV}/c$ beam momentum, recorded in 2022). Fig.5 bottom right panel shows the same high-multiplicity data set with only a single Sense Wire switched on, thus eliminating the cartesian combinatorial ghost hits using the design feature explained in Section 3.
@@
Figure 5: (Color online) Top panel: an example for the $\Delta{y}$ versus $y$ drift velocity calibration method on a real data sample in a $5$-minute time window. The corresponding data set is a low multiplicity run (proton-carbon collisions at $120\,\mathrm{GeV}/c$ beam momentum, recorded in 2023). Here, the full GRC acceptance was used. Bottom left panel: the equivalent plot for a high multiplicity run, with the full GRC acceptance (minimum-bias Pb+Pb collisions at $150\,A\mathrm{GeV}/c$ beam momentum, recorded in 2022). Bottom right panel: the same high multiplicity data set with only a single Sense Wire used, thus eliminating the cartesian combinatorial ghost hits in GRC. Note the substantially-reduced background. The color scale is merely to emphasize hit density.
@@
In order to further suppress the combinatorial background accompanying the correlation signalsignal, the ROOT implementation of the Least Trimmed Squares (LTS) fitting was used. After calibrating MTPC-L, the upstream chambers are subsequently calibrated step-by-step using downstream chambers as geometric references.
@@
The statistical uncertainty of the method can be estimated from above by treating the pickup wire electrodes as $6\,\mathrm{mm}$ wide strips. Thus, the single-track position resolution of the GRC is not worse than $6\,\mathrm{mm}/\sqrt{12}$, corresponding to the resolution of a $6\,\mathrm{mm}$ wide uniform distribution. For a track sample of a ${\sim}5\,\mathrm{minute}$ time window, typically there are hundreds of MTPC-L tracks hitting the GRC. It is thus possible to reach the desired statistical uncertainty of $1\,\mathrm{mm}$, corresponding to $1$ per mil of the ${\sim}1\,\mathrm{m}$ TPC drift length. The systematic accuracy of the method was quantified using a closure test. By re-reconstructing the data using the calibrated drift velocities, the calculated correction on top of the previously-corrected drift velocities was found to be better than 1-2 permil. This can be attributed to the systematic uncertainty of the method. The systematics were also double-checked by comparing the obtained drift velocities to the exhaust analysis method, which agreed within 1-2 permil, up to a normalization constant. Note that the exhaust analysis method is expected to have worse systematics in comparison to the more-direct $\Delta{y}$ versus $y$ method discussed in this paper. These validation plots are seen in Fig.6.
@@
Figure 6: (Color online) Final validation of the GRC-based $\Delta{y}$ versus $y$ drift velocity calibration method. It is seen that the closure test is satisfied to the desired accuracy (1-2 permil). Moreover, good consistency is seen with the exhaust analyzer method. (Note that the exhaust analyzer method is expected to have some absolute systematic error, hence the necessity for its normalization scale adjustment.)
@@ 5 Concluding remarks
In this paper a novel cost-efficient method was described for in-situ drift velocity monitoring in large volume TPCs. The method was deployed and is currently in use at the NA61/SHINE experiment at CERN. The key idea is to place a low-cost segmented detector, called to be the Geometry Reference Chamber (GRC), downstream of the TPC to be monitored. This monitoring solution was added retrospectively to the existing TPC system without a surgical operation to the chambers. A GRC was cost-effectively realized by using a MWPC designed to be compatible with existing TPC readout electronics. In such a way the development time and cost of the readout electronics could be spared. In order to match such an MWPC to the TPC readout electronics, it was necessary to slow down the signal formation in the GRC, since TPC readout electronics typically posses a considerable trigger latency. This signal slowing was achieved by using an adjustable drift field transverse to the particle passage direction. The number of required readout channels was minimized using cartesian readout. Since the NA61/SHINE experiment is used to study both low- and high-multiplicity collisions, the acceptance of the GRC was designed to be sufficiently large while addressing the issue of combinatorial ghost hits. The latter was handled by making the acceptance adjustable: when studying high-multiplicity collisions, individual Sense Wires can be switched off down to a single wire, eliminating the cartesian ghosts, whereas for low-multiplicity runs the full acceptance is used with all of the Sense Wires operational.
@@
Motivated by the GRC-based drift velocity monitoring method, a calibration procedure for the other geometric calibration parameters of the NA61/SHINE TPC chambers was also developed. Using that concept, all 8 parameters (time-dependent drift velocity, trigger latency, alignment shifts, and alignment angles) of the TPC system were successfully calibrated.
@@ Acknowledgements
Appendix A Trigger latency and drift-direction alignment of TPCs
@@
as seen from Eq.(4.4). Taking several data samples with in-situ drift velocity differing by $1-2\%$, the slope of the $\Delta{y}$ versus $v_{\mathrm{drift}}$ scatter plot provides the additive correction between the true trigger latency ($t_{0,\mathrm{true}}$) against the trigger latency estimate applied during the reconstruction ($t_{0,\mathrm{nom}}$). Thus, $t_{0,\mathrm{true}}$ can be determined from the TPC-GRC mismatch data. After the $t_{0}$ is calibrated, the TPC-GRC mismatch will be not only $y$-independent, but also $v_{\mathrm{drift}}$-independent, and one is left with $\Delta{y}=y_{\mathrm{anode,nom}}-y_{\mathrm{anode,true}}$. Thus, the drift direction displacement ($y$-alignment correction) $y_{0}:=y_{\mathrm{anode,nom}}-y_{\mathrm{anode,true}}$ of the TPC versus the GRC can be determined from the TPC-GRC mismatch data. An example for the $\Delta{y}$ versus $v_{\mathrm{drift}}$ calibration method is shown in Fig.7. Having the $t_{0}$ and $y_{0}$ obtained, the upstream TPC chambers can also be calibrated, using the already calibrated TPCs as geometry reference, in place of the GRC.
Figure 7: (Color online) An example for the $\Delta{y}$ versus $v_{\mathrm{drift}}$ analysis for trigger latency ($t_{0}$) and drift direction shift ($y_{0}$) calibration, after the drift velocity ($v_{\mathrm{drift}}$) has been obtained. The slope of the pertinent scatter data gives the $t_{0}$ correction, whereas the offset gives the $y_{0}$ correction. For the method to be workable, data samples with different drift velocities (up to $1-2\%$) are used. The pertinent data set is a high multiplicity collision type (minimum-bias Pb+Pb collisions at $150\,A\mathrm{GeV}/c$ beam momentum, recorded in 2022).
@@ Appendix B Systematic errors and alignment calibration of the TPCs
As summarized in Section 4, the final systematics of the in-situ drift velocity calibration via the $\Delta{y}$ versus $y$ method was verified to be not worse than 1-2 permil (or $1\,\mathrm{mm}$ overall, along the ${\sim}1\,\mathrm{m}$ drift length). This performance, however, was only achievable after some detailed studies. The main systematics error comes from a possible misalignment of the adjacent chambers, or misalignment of MTPC-L versus the GRC. Assume that we have a pair of TPC chambers, to be calibrated against each-other from the point of view of drift velocity. Assume moreover that the downstream one is either the GRC, or is an already drift velocity calibrated TPC. Using automatic formula manipulation programs (e.g. Maple) evaluating affine transformations is not difficult, and it is rather straightforward to show that if the upstream chamber has an imperfectly known alignment, that contributes to the multiplicative drift velocity correction $v_{\mathrm{drift,corr}}:=v_{\mathrm{drift,true}}/v_{\mathrm{drift,nom}}$ obtained by the $\Delta{y}$ versus $y$ in-situ drift velocity calibration method, described in Section 4. The track-by-track systematics is seen to be
@@
$\mathrm{syst}(v_{\mathrm{drift,corr}})=\Big{(}y\cdot\theta_{x}+x\cdot\theta_{y}+(y-y_{\mathrm{mainvertex}})\cdot\theta_{x}+z_{0}\Big{)}\cdot\frac{1}{z-z_{\mathrm{mainvertex}}}$ (B.1)
@@
to the first order. [^5] This can be derived analytically, for magnetic field-off data, i.e. for straight tracks. Since the drift direction of the NA61/SHINE TPCs approximately coincide with the direction of the magnetic bending field, i.e. the drift and the bending does not interfere to the first order, the formula holds for magnetic field-on data as well, to a good approximation. Here, $x,y,z$ is the particle hit position at the $z=const$ reference plane (e.g. the GRC plane) of the $\Delta{y}$ versus $y$ study, moreover $x_{0},y_{0},z_{0}$ denote the (unkown) correction to the position alignment of the upstream chamber, and $\theta_{x},\theta_{y},\theta_{z}$ denote the (unknown) correction to the angular alignment of the upstream chamber, whereas the $x,y,z$ coordinates indexed by $()_{\mathrm{mainvertex}}$ denote the coordinates of the collision point. [^6] The downstream chamber is assumed to be either the alignment reference or to be already calibrated for alignment. In our experiment, MTPC-L is taken as reference for alignment, and all the other chambers are calibrated against this reference, pairwise. It is seen that the drift velocity systematics decreases with distance from the main-vertex, and therefore justifies our placement of the GRC downstream of all the TPCs, as seen in Fig.2. Even if the angular or position alignment of the MTPC-L with respect to the GRC were imperfectly known to $|y\cdot\theta_{x}|\lesssim 1\,\mathrm{cm}$ or $|x\cdot\theta_{y}|\lesssim 1\,\mathrm{cm}$ or $|(y-y_{\mathrm{mainvertex}})\cdot\theta_{x}|\lesssim 1\,\mathrm{cm}$ or $|z_{0}|\lesssim 1\,\mathrm{cm}$, since $|z_{{}_{\mathrm{GRC}}}-z_{\mathrm{mainvertex}}|\approx 16\,\mathrm{m}$, this would not disturb the MTPC-L versus GRC drift velocity calibration, beyond half a permil. That is, the alignment imperfection does not give sizable systematics to the MTPC-L versus GRC in-situ drift velocity calibration. When propagating the drift velocity calibration toward the upstream chambers (VTPC-1, VTPC-2), however, the systematics Eq.(B.1) by imperfectly known alignment can give a sizable contribution. Therefore, in this appendix we briefly describe, how the alignment is calibrated.
@@
A TPC chamber has in total 8 unknown calibration coefficients related to its geometry: the $x_{0},y_{0},z_{0}$ position misalignment correction (transverse, drift direction, longitudinal), the $\theta_{x},\theta_{y},\theta_{z}$ angular misalignment correction, the drift velocity correction $(v_{\mathrm{drift,corr}}-1)$, and the correction to the trigger latency $t_{0,\mathrm{corr}}:=t_{0,\mathrm{true}}-t_{0,\mathrm{nom}}$. These unkown imperfection parameters, to be determined by the calibration procedure, are small, and therefore, correction to the first order Taylor expansion terms with respect to these is enough. We use special magnetic field-off calibration runs in order to have only straight trajectories for charged particle tracks throughout the chamber system. In other words, the chamber system is tomographed by an ensemble of straight global tracks associated to the main-vertex, and their mismatch pattern is studied when dissected and refitted by straight lines in the local detectors, as illustrated in Fig.8.
@@
Figure 8: (Color online) Sketch of the alignment calibration strategy for adjacent TPCs: special magnetic field-off alignment calibration runs are recorded, i.e. the chamber system is tomographed by straight global main-vertex tracks. If the relative alignment parameters for the chambers assumed during the reconstruction are imperfect, the track parameters will have a mismatch, when compared on a joint reference plane, for the adjacent chambers.
@@
Quantitatively, the magnetic field-off data based alignment calibration goes as follows. For MTPC-L, the $(v_{\mathrm{drift,corr}}-1)$ parameter is determined as in Section 4, and the parameters $t_{0,\mathrm{corr}}$,$y_{0}$ are determined as in Appendix A, whereas $\theta_{x}=\theta_{y}=\theta_{z}=0$ and $x_{0}=z_{0}=0$ by convention, i.e. MTPC-L is taken as alignment reference. The other chambers are calibrated against each-other pairwise, for all the 8 parameters, assuming that e.g. the downstream one is already calibrated. An imperfection in the assumption about the 8 parameters of the upstream chamber causes a mismatch between locally refitted track pieces in the adjacent chambers, when compared on a $z=const$ reference plane, see again Fig.8. If the straight track segments are parameterized by the four parameters $N_{x},N_{y},M_{x},M_{y}$ as $x=M_{x}+(z-z_{\mathrm{ref}})\cdot N_{x}$ and $y=M_{y}+(z-z_{\mathrm{ref}})\cdot N_{y}$, then the locally refitted and compared local tracks will give a track parameter mismatch vector field $\Delta{N}_{x},\Delta{N}_{y},\Delta{M}_{x},\Delta{M}_{y}$ as a function of the track parameters $N_{x},N_{y},M_{x},M_{y}$ in the downstream chamber (which is already calibrated or reference). Evaluating affine transformations, one arrives at the formula
@@
$\Delta{M}_{x}$ $=$ $N_{x}\cdot(M_{y}\cdot\theta_{x}+M_{x}\cdot\theta_{y}+z_{0})+z\cdot\theta_{y}+M_{y}\cdot\theta_{z}-x_{0},$ (B.4)
$\Delta{M}_{y}$ $=$ $N_{y}\cdot(M_{x}\cdot\theta_{y}+M_{y}\cdot\theta_{x}+z_{0})+z\cdot\theta_{x}-M_{x}\cdot\theta_{z}-y_{0}$ (B.6)
@@
Figure 9: (Color online) Top panels: an example for the $\theta_{x},\theta_{y},\theta_{z}$ angular alignment calibration using $\Delta{N}_{x}$ versus $N_{x}$ and $\Delta{N}_{y}$ versus $N_{x}$ plots for $|N_{y}|\approx 0$ tracks in magnetic field-off data (the pertinent sample is minimum-bias Pb+Pb at $150\,A\mathrm{GeV}/c$ beam momentum, recorded in 2022). Bottom panel: an example for the $x_{0},z_{0}$ position alignment calibration using $\Delta{M}_{x}$ versus $N_{x}$ plot for tracks in magnetic field-off data (using the same sample as for the top panels). The black vertical lines indicate fiducial cuts, in order to avoid regions close to the TPC field cages, where the drift electric field can be imperfect (inhomogeneus). The color scale is merely for emphasizing hit density.

### v2

@@ 4 Drift velocity calibration using the GRC
Here $y_{\mathrm{nom}}$ is the reconstructed drift coordinate assuming a nominal estimate during reconstruction for the position of the TPC amplification plane at $y_{\mathrm{anode},\mathrm{nom}}$, the anode meaning the field wire / sense wire plane of the TPC. The symbol $t_{0,\mathrm{nom}}$ stands for a nominal estimate for the trigger delay during reconstruction. $t_{\mathrm{drift,raw}}$ is the measured particle track cluster position in terms of drift time as measured by the TPC readout. $v_{\mathrm{drift,nom}}$ is the nominal estimate for the electron drift velocity in the TPC during reconstruction. The quantities labeled by $\,{()}_{\mathrm{true}}\,$ are the true but unknown values of these calibration factors. The negative sign in Eq.(4.2) is due to the coordinate convention in NA61/SHINE. The coordinate $y_{\mathrm{nom}}$ measured by the TPC along the drift direction corresponds to a GRC measurement $y_{\mathrm{true}}$, for each track matched to a GRC hit, as seen in Fig.1. The matching is done via a typically ${\sim}3\,\mathrm{cm}$ wide tolerance window. By eliminating the running variable $t_{\mathrm{drift,raw}}$ in Eq.(4.2), one infers
@@
for the ensemble of TPC-GRC tracks. Here $\Delta{y}:=y_{\mathrm{nom}}-y_{\mathrm{true}}$ stands for the drift coordinate TPC-GRC mismatch and $y_{\mathrm{true}}$ stands for the GRC measurement. Occasionally, the label $()_{\mathrm{true}}$ will be suppressed where not confusing. The slope of the $\Delta{y}$ versus $y$ scatter plot determines the correction factor between the true drift velocity $v_{\mathrm{drift,true}}$ and the estimate $v_{\mathrm{drift,nom}}$ assumed during the reconstruction, thus $v_{\mathrm{drift,true}}$ can be estimated from the TPC-GRC mismatch data. This is shown in Fig.5 top panel for a typical low-multiplicity data set (proton-carbon collisions at $120\,\mathrm{GeV}/c$ beam momentum, recorded in 2023). In Fig.5 bottom left panel the same is shown, but for a typical high-multiplicity data set (Pb+Pb collisions at $150\,A\mathrm{GeV}/c$ beam momentum, recorded in 2022). Fig.5 bottom right panel shows the same high-multiplicity data set with only a single sense wire switched on, thus eliminating the Cartesian combinatorial ghost hits using the design feature explained in Section 3.
@@
Figure 5: (Color online) Top panel: an example for the $\Delta{y}$ versus $y$ drift velocity calibration method on a real data sample in a $5$-minute time window. The corresponding data set is a low multiplicity run (proton-carbon collisions at $120\,\mathrm{GeV}/c$ beam momentum, recorded in 2023). Here, the full GRC acceptance was used. Bottom left panel: the equivalent plot for a high multiplicity run, with the full GRC acceptance (Pb+Pb collisions at $150\,A\mathrm{GeV}/c$ beam momentum, recorded in 2022). Bottom right panel: the same high multiplicity data set with only a single sense wire used, thus eliminating the Cartesian combinatorial ghost hits in GRC. Note the substantially reduced background. The color scale of the scatter plot is merely to emphasize hit density for the eye. The line on the plots corresponds to a straight line fit to the scatter data, using the background tolerant LTS method. The slopes of the fitted lines yield the drift velocity correction.
@@
In order to further suppress the combinatorial background accompanying the correlation signal, the ROOT [18, 19] implementation of the Least Trimmed Squares (LTS) fitting was used. After calibrating MTPC-L, the other chambers in the TPC system are subsequently calibrated step-by-step against each other, using the already calibrated ones as geometric reference.
@@
The statistical uncertainty of the method can be estimated from above by treating the pickup wire electrodes as $6\,\mathrm{mm}$ wide strips. Thus, the single-track position resolution of the GRC is not worse than $6\,\mathrm{mm}/\sqrt{12}$, corresponding to the resolution of a $6\,\mathrm{mm}$ wide uniform distribution. For a track sample of a ${\sim}5\,\mathrm{minute}$ time window, typically there are hundreds of MTPC-L tracks hitting the GRC, giving a resolution of $6\,\mathrm{mm}/\sqrt{12}/\sqrt{N_{\mathrm{tracks}}}$ for such ensemble of tracks. It is thus possible to push the statistical uncertainty below the desired $1\,\mathrm{mm}$, corresponding to $1$ per mil of the ${\sim}1\,\mathrm{m}$ TPC drift length. The systematic accuracy of the method was quantified using a closure test. By re-reconstructing the data using the calibrated drift velocities, the calculated correction on top of the previously-corrected drift velocities was found to be better than 1-2 permil. The systematics were also double-checked by comparing the obtained drift velocities to the exhaust analysis method, which agreed within 1-2 permil, up to a scale adjustment constant. Note that the exhaust analysis method is expected to have worse systematic uncertainties, hence the need for the scale adjustment constant for the exhaust method. These validation plots are seen in Fig.6.
@@
Figure 6: (Color online) Top panel: final validation of the GRC-based $\Delta{y}$ versus $y$ drift velocity calibration method. Statistical uncertainties of the data points are below $0.05\%$, therefore errorbars are not shown. It is seen that the closure test is satisfied to the desired accuracy (1-2 permil). That provides an empirical estimate for the systematic uncertainties. Moreover, consistency is seen with the exhaust analyzer method, up to a scale adjustment factor of $1.0132$. Note that the exhaust analyzer method is expected to have some absolute systematic error, which explains the necessity for the empirical scale adjustment factor. Bottom panel: the same data points, shown in terms of relative difference with respect to the GRC-based calibration method.
After calibrating the MTPC-L using the GRC, the other chambers are calibrated pairwise against each-other. Namely, the calibration scheme is MTPC-L from GRC, followed by VTPC-2 from MTPC-L, then MTPC-R from VTPC-2, and VTPC-1 from VTPC-2, and so on. Once having a methodology for drift velocity calibration, it is possible to extract further calibration constants from the data. Namely, it is possible to estimate the the trigger delay ($t_{0}$) and the drift direction displacement ($y_{0}$) of a chamber, see Appendix A. Moreover, using magnetic field-off calibration data, it is possible to estimate the chamber alignment parameters, as explained in Appendix B. A detailed systematic error analysis is also provided there. It is worth to note that after the above drift velocity and geometry calibration procedure, the magnetic field-off data are also used to map out the electric drift field distortions inside the TPCs: residuals of refitted local tracks are tabulated as a function of space, and are used as a fine-correction. At the boundaries, these residuals are up to ${\sim}1\,\mathrm{mm}$ and are much smaller inside the volumes. The drift field distortion corrections are of second order, and therefore do not interfere with the discussed drift velocity and geometry calibration.
@@ 5 Concluding remarks
In this paper a novel cost-efficient method was described for in-situ drift velocity monitoring in large volume TPCs. The method was deployed and is currently in use at the NA61/SHINE experiment at CERN. The key idea is to place a low-cost segmented detector, called to be the Geometry Reference Chamber (GRC), downstream of the TPC to be monitored. This monitoring solution was added retrospectively to the existing TPC system without a surgical operation to the chambers. A GRC was cost-effectively realized by using a MWPC designed to be compatible with existing TPC readout electronics. In such a way the development time and cost of the readout electronics could be spared. In order to match such an MWPC to the TPC readout electronics, it was necessary to slow down the signal formation in the GRC, since TPC readout electronics typically allows for a considerable trigger delay. This signal slowing was achieved by using an adjustable drift field transverse to the particle passage direction. The number of required readout channels was minimized using Cartesian readout. Since the NA61/SHINE experiment is used to study both low- and high-multiplicity collisions, the acceptance of the GRC was designed to be sufficiently large while addressing the issue of combinatorial ghost hits. The latter was handled by making the acceptance adjustable: when studying high-multiplicity collisions, individual sense wires can be switched off down to a single wire, eliminating the Cartesian ghosts, whereas for low-multiplicity runs the full acceptance is used with all of the sense wires operational.
@@
Motivated by the GRC-based drift velocity monitoring method, a calibration procedure for the other geometric calibration parameters of the NA61/SHINE TPC chambers was also developed. Using that concept, all 8 parameters (time-dependent drift velocity, trigger delay, alignment shifts, and alignment angles) of the TPC system were successfully calibrated, for each chamber.
@@ Acknowledgements
Appendix A Trigger delay and drift-direction alignment of TPCs
@@
as seen from Eq.(4.4). Taking several data samples with actual drift velocity differing by $2-3\%$, the slope of the $\Delta{y}$ versus $v_{\mathrm{drift}}$ scatter plot provides an estimate for the additive correction between the true trigger delay ($t_{0,\mathrm{true}}$) and the trigger delay rough estimate applied during the initial reconstruction ($t_{0,\mathrm{nom}}$). Thus, $t_{0,\mathrm{true}}$ can be estimated from the TPC-GRC mismatch data. After the $t_{0}$ is calibrated, the TPC-GRC mismatch will be not only $y$-independent, but also $v_{\mathrm{drift}}$-independent, and one is left with $\Delta{y}=y_{\mathrm{anode,nom}}-y_{\mathrm{anode,true}}$. Thus, the drift direction displacement ($y$-alignment correction) $y_{0}:=y_{\mathrm{anode,nom}}-y_{\mathrm{anode,true}}$ of the TPC versus the GRC can be estimated from the TPC-GRC mismatch data. An example for the $\Delta{y}$ versus $v_{\mathrm{drift}}$ calibration method is shown in Fig.7. Having the $t_{0}$ and $y_{0}$ obtained, the upstream TPC chambers can also be calibrated, using the already calibrated TPCs as geometry reference, in place of the GRC.
Figure 7: (Color online) An example for the $\Delta{y}$ versus $v_{\mathrm{drift}}$ analysis for trigger delay ($t_{0}$) and drift direction shift ($y_{0}$) calibration, after the drift velocity ($v_{\mathrm{drift}}$) has been obtained. The slope of the pertinent scatter data gives the $t_{0}$ correction, whereas the offset gives the $y_{0}$ correction. For the method to be workable, data samples with different drift velocities (up to $2-3\%$) are used. The pertinent data set is a high multiplicity collision type (Pb+Pb collisions at $150\,A\mathrm{GeV}/c$ beam momentum, recorded in 2022).
@@ Appendix B Systematic errors and alignment calibration of the TPCs
As summarized in Section 4, the final systematic errors of the in-situ drift velocity calibration via the $\Delta{y}$ versus $y$ method was verified to be not worse than 1-2 permil (or $1\,\mathrm{mm}$ overall, along the ${\sim}1\,\mathrm{m}$ drift length). This performance, however, was achievable after some detailed studies. Since the adjacent TPCs are calibrated against each-other (with first MTPC-L being calibrated against the GRC), the main source of systematic errors is the possible misalignment of the adjacent chambers (or the misalignment of MTPC-L versus the GRC). Considering a pair of adjacent chambers, one of which is an already fully calibrated chamber or the GRC, the other being an uncalibrated TPC, evaluation of affine transformations show that the track-by-track systematic error in the multiplicative drift velocity correction $v_{\mathrm{drift,corr}}:=v_{\mathrm{drift,true}}/v_{\mathrm{drift,nom}}$ caused by misalignment is
@@
$\mathrm{syst}(v_{\mathrm{drift,corr}})=\Big{(}y\cdot\theta_{x}+x\cdot\theta_{y}+(y-y_{\mathrm{mainvertex}})\cdot\theta_{x}+z_{0}\Big{)}\cdot\frac{1}{z-z_{\mathrm{mainvertex}}}\,,$ (B.1)
@@
to the first order. [^5] This identity can be derived analytically, for magnetic field-off data, i.e. for straight tracks. Since the drift directions of the NA61/SHINE TPCs approximately coincide with the direction of the magnetic bending field, i.e. the drift and the bending does not interfere to the first order, the formula holds for magnetic field-on data as well, to a good approximation. In order to aid the tedious analytic evaluation of the three dimensional affine transformations, automatic formula manipulation software was used (the LinearAlgebra package of Maple). Here, $x,y,z$ is the particle hit position at the $z=const$ reference plane of the $\Delta{y}$ versus $y$ study (e.g. the GRC plane), moreover $x_{0},y_{0},z_{0}$ denote the (unknown) correction to the position alignment of the uncalibrated chamber, and $\theta_{x},\theta_{y},\theta_{z}$ denote the (unknown) correction to the angular alignment of the uncalibrated chamber, whereas the $x,y,z$ coordinates indexed by $()_{\mathrm{mainvertex}}$ denote the coordinates of the collision point. It is seen that the alignment caused drift velocity systematic error decreases with distance from the main-vertex, and therefore justifies our placement of the GRC downstream of all the TPCs, as seen in Fig.2. Even if the alignment of the MTPC-L with respect to the GRC were imperfectly known to $|y\cdot\theta_{x}|\lesssim 1\,\mathrm{cm}$ or $|x\cdot\theta_{y}|\lesssim 1\,\mathrm{cm}$ or $|(y-y_{\mathrm{mainvertex}})\cdot\theta_{x}|\lesssim 1\,\mathrm{cm}$ or $|z_{0}|\lesssim 1\,\mathrm{cm}$, this would not disturb the MTPC-L versus GRC drift velocity calibration beyond half permil, since $|z_{{}_{\mathrm{GRC}}}-z_{\mathrm{mainvertex}}|\approx 16\,\mathrm{m}$ is large. That is, the alignment imperfection does not give sizable systematic error to the MTPC-L versus GRC drift velocity calibration. When propagating the drift velocity calibration toward the upstream chambers (VTPC-1, VTPC-2), however, the systematic error Eq.(B.1) by imperfectly known alignment can give a sizable contribution. Therefore, in this appendix we briefly describe, how the alignment is calibrated.
@@
A TPC chamber has in total 8 unknown calibration coefficients related to its geometry: the $x_{0},y_{0},z_{0}$ position misalignment correction (transverse, drift direction, longitudinal), the $\theta_{x},\theta_{y},\theta_{z}$ angular misalignment correction, the drift velocity correction $(v_{\mathrm{drift,corr}}-1)$, and the correction to the trigger delay $t_{0,\mathrm{corr}}:=t_{0,\mathrm{true}}-t_{0,\mathrm{nom}}$. These unknown imperfection parameters, to be determined by the calibration procedure, are small, and therefore correction to the first order Taylor expansion terms with respect to these is enough. We use special magnetic field-off calibration runs in order to have only straight trajectories for charged particle tracks throughout the chamber system. In other words, a tomography of the chamber system is done using an ensemble of straight global tracks associated to the main-vertex, and their mismatch pattern is studied when dissected and refitted by straight lines in the local detectors, as illustrated in Fig.8.
@@
Figure 8: (Color online) Sketch of the alignment calibration strategy for adjacent TPCs: special magnetic field-off alignment calibration runs are recorded, i.e. a tomography of the chamber system is performed by straight global main-vertex tracks. If the relative alignment parameters for the chambers assumed during the reconstruction are imperfect, the track parameters will have a mismatch, when compared on a joint reference plane, for the adjacent chambers.
@@
Quantitatively, the magnetic field-off data based alignment calibration goes as follows. For MTPC-L, the $(v_{\mathrm{drift,corr}}-1)$ parameter is determined as in Section 4, and the parameters $t_{0,\mathrm{corr}}$,$y_{0}$ are determined as in Appendix A, whereas $\theta_{x}=\theta_{y}=\theta_{z}=0$ and $x_{0}=z_{0}=0$ by convention, i.e. MTPC-L is taken as alignment reference. The other chambers are calibrated against each-other pairwise, for all the 8 parameters, assuming that one in the pair is already calibrated. In order to fix sign conventions, in this section it is assumed that the downstream chamber is calibrated, but analogous calibration equations hold when the upstream chamber is the calibrated one. An imperfection in the assumption about the 8 parameters of the upstream chamber causes a mismatch between locally refitted track pieces in the adjacent chambers, when compared on a $z=const$ reference plane, see again Fig.8. If the straight track segments are parameterized by the four parameters $N_{x},N_{y},M_{x},M_{y}$ as $x(z)=M_{x}+(z-z_{\mathrm{ref}})\cdot N_{x}$ and $y(z)=M_{y}+(z-z_{\mathrm{ref}})\cdot N_{y}$, then the locally refitted and compared local tracks will give a track parameter mismatch vector field $\Delta{N}_{x},\Delta{N}_{y},\Delta{M}_{x},\Delta{M}_{y}$ as a function of the track parameters $N_{x},N_{y},M_{x},M_{y}$ in the downstream chamber (which is already calibrated or reference). Evaluating affine transformations, one arrives at the formula
@@
$\Delta{M}_{x}$ $=$ $N_{x}\cdot(M_{y}\cdot\theta_{x}+M_{x}\cdot\theta_{y}+z_{0})+z_{\mathrm{ref}}\cdot\theta_{y}+M_{y}\cdot\theta_{z}-x_{0},$ (B.4)
$\Delta{M}_{y}$ $=$ $N_{y}\cdot(M_{x}\cdot\theta_{y}+M_{y}\cdot\theta_{x}+z_{0})+z_{\mathrm{ref}}\cdot\theta_{x}-M_{x}\cdot\theta_{z}-y_{0}$ (B.6)
@@
Figure 9: (Color online) Top panels: an example for the $\theta_{x},\theta_{y},\theta_{z}$ angular alignment calibration using $\Delta{N}_{x}$ versus $N_{x}$ and $\Delta{N}_{y}$ versus $N_{x}$ plots for $|N_{y}|\approx 0$ tracks in magnetic field-off data (the pertinent sample is Pb+Pb at $150\,A\mathrm{GeV}/c$ beam momentum, recorded in 2022). Bottom panel: an example for the $x_{0},z_{0}$ position alignment calibration using $\Delta{M}_{x}$ versus $N_{x}$ plot for tracks in magnetic field-off data (using the same sample as for the top panels). The black vertical lines indicate fiducial cuts, in order to avoid regions close to the TPC field cages, where the drift electric field can be imperfect (inhomogeneous). The color scale of the scatter plots is merely for emphasizing hit density for the eye. The line on the plots corresponds to a straight line fit to the scatter data, using the background tolerant LTS method.